OpenStudy (anonymous):

A machine contains 125 gumballs. If the probability of getting a red gumball is 3/5. What is the probability of NOT getting a red gumball?

OpenStudy (anonymous):

Hmm,what do u think? @FRUT

OpenStudy (mrnood):

it doesn't matter how many sweets there are the prob of something NOT happening is (1- probabilty of happening)

OpenStudy (anonymous):

\[1-\frac{ 3 }{ 5 }=\frac{ 5 }{ 5 }-\frac{ 3 }{ 5 }\]\[\frac{ 5-3 }{ 5 }=\frac{ ? }{ ? }\]solve it @FRUT

OpenStudy (anonymous):

2/5